OpenStudy (anonymous):

please helppp..What is the value of a for the following circle in general form? x^2+y^2+ax+by+c=0

OpenStudy (anonymous):

you will need to complete the square for both x and y general equation: \[x^{2} + ax = (x + a/2)^{2} - (a/2)^2\]

OpenStudy (anonymous):

help...

OpenStudy (anonymous):

lol ok. well that was for x. \[y^{2} + bx = (y + b/2)^{2} - (b/2)^{2}\] completing the square just gives an equivalent way of writing it. and the point of it is to put it in the general form of the circle.

OpenStudy (anonymous):

so: \[(x +a/2)^{2} + (y + b/2)^{2} = (a/2)^{2} + (b/2)^{2}\]

OpenStudy (anonymous):

yeah i think

OpenStudy (anonymous):

r ^ 2 = [(a/2)^2 + (b/2)^2]

OpenStudy (anonymous):

so how would i find the value of a

OpenStudy (anonymous):

the final answer is 3 above

OpenStudy (anonymous):

you can't, it's not given. the question was just to place it in the general form

OpenStudy (anonymous):

my bad . i didn't see the c

OpenStudy (anonymous):

so is the question like a bad question,it's asking for the value of a

OpenStudy (anonymous):

final answer is : \[(x + a/2)^2 + (y + b/2)^2 = [(a/2)^2 + (b/2)^2 - c]\]

OpenStudy (anonymous):

ugh this is so confusingggggggggggg i dont know what to type in

OpenStudy (anonymous):

i'm not sure what they mean by value of a. maybe by a they mean h. and the answer would be -a/2

OpenStudy (anonymous):

or maybe a is the radius

OpenStudy (anonymous):

i have no idea,i copy and pasted the equation

OpenStudy (anonymous):

do they show this somewhere? \[(x - a)^2 + (y - b)^2 =r^2\] or do they use h and k ?

OpenStudy (anonymous):

at the very top is all they show and the graph thaty's it]